Suppose prices in the U.S. increase. If euro prices are constant and with no change in...

  1. Suppose prices in the U.S. increase. If euro prices are constant and with no change in the nominal exchange rate, the euro has undergone:
    1. Real appreciation
    2. Real depreciation
    3. Nominal appreciation
    4. Nominal depreciation
    5. None of these
  2. According to PPP, what is the long-run relationship between inflation and the nominal exchange rate:
    1. Higher inflation leads to depreciation
    2. Ambiguous relationship-it depends on extend of the Fed’s response
    3. Higher inflation leads to appreciation
Answer #1
  1. The exchange rate has important relationship to the price level because it represents a link between domestic prices and foreign prices.

Here we can take and example to understand:

Let the price in US be $100

Exchange rate = $1.5/£

Then the price in Euro= 100/1.5= 66.67

Now

If the domestic price increase from 100 to 110, the price in EURo will be :

110= 1.5 * x

X= 110/1.5= 73.33

So the price has increased, while the exchange rate remain same . This situation is called real depreciation.

So the ans. Is B

2.

Relative PPP says that the changes in exchange rate between two countries is equals to the difference between the inflation rates between the two countries. And higher the inflation , it will lead to increase in the absolute value of currency, but decrease the actual value, so higher inflation leads to depreciation.

So the answer is A.